Use the vertex form of the quadratic equation to find the vertex.
y = (x-2)^2 + 4

Answer by stanbon(57246)

Use the vertex form of the quadratic equation to find the vertex.
y = (x-2)^2 + 4
(x-2)^2 = y-4
----
Vertex: (2,4)
=================
Cheers,
Stan H.